OpenStudy (anonymous):

will someone kindly help me get the correct answer,,,, raise to the power indicated: (a^2bc^3)^3

hartnn (hartnn):

just 2 simple rules : \(\large (xy)^m=x^my^m \\ \large (x^m)^n=x^{mn }\)

OpenStudy (anonymous):

(xy)^m=x^my^m